How to solve something that has values a=0.05 b=0 and c=6, i need help ASAP
Found 2 solutions by TimothyLamb, DrBeeee:
Answer by TimothyLamb(4379)

a=0.05 b=0 and c=6 are the coefficients of a quadratic equation:
0.05x^2 + 0x + 6 = 0
---
the above quadratic equation is in standard form, with a=0.05, b=0, and c=6
---
to solve the quadratic equation, plug this:
0.05 0 6
into this: https://sooeet.com/math/quadratic-equation-solver.php
---
the roots (the solutions) of this quadratic equation are complex:
0 + (10.9544512)i
0 - (10.9544512)i
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php

Answer by DrBeeee(684)

We are given the values of three variables that make a function of these three variables equal to zero. These values are called zeroes (or roots in my day). What we need to find the factor that has that zero. This done by simply eqatipng the variable to the zero, then get the factor by subtracting the zero from each side. Confused? I'm not very good at explanations, but I'm sure you'll see what I mean with the "arithmetic".
Let's start with the variable a. It has the zero at
(1) a = 0.05
Now subtract 0.05 from each side of (1) to get
(2) a - 0.05 = 0.05 - 0.05 or
(3) a - 0.05 = 0
The expression (3) is your first factor
(4) (a-0.05)
Next set
(5) b = 0 or
(6) b - 0 = 0 so our second factor is just
(7) (b)
Now set the third variable to its zero to get
(8) c = 6
Subtract 6 from each side to get
(9) c - 6 = 6 - 6 or
(10) c - 6 = 0 so our third factor is
(11) (c-6)
Combine the three factors (by multiplying) to form the function
(12) f(a,b,c) = (a-0.05)*(b)*(c-6)
Answer: The "something" you solve is (a-0.05)*b*(c-6) = 0